# Unifying the Perspectives of NLP and Software Engineering: A Survey on Language Models for Code

### v3

@@ 3.1.1 Text-to-Code
Text-to-code tasks take text as input, and output code. Related methods are listed in Figure 4
@@ 3.1.2 Code-to-Code
Code-to-code tasks take code as input, and output code. Related methods are listed in Figure 5 and 6
@@
[Code Evaluation
[Text-to-Code
[Code-to-Code
[Code Search
@@
COSAL (Mathew & Stolee, 2021)
Corder (Bui et al., 2021b), Cosco (Saieva et al., 2023)
@@
Completion
@@
Deep3 (Raychev et al., 2016a), Hellendoorn & Devanbu (2017)
@@
SLM (Alon et al., 2020), Pythia (Svyatkovskiy et al., 2019), NLM (Karampatsis et al., 2020)
@@
LongCoder (Guo et al., 2023b)
@@
Translation
@@
mppSMT (Nguyen et al., 2015)
Tree2Tree (Chen et al., 2018), Grammar Tree2Tree (Drissi et al., 2018)
@@
Self-Debugging (Chen et al., 2023c),
@@
Pan et al. (2023), Tang et al. (2023)
@@
Mining
RASH (Zhang et al., 2018), RAPIM (Alrubaye et al., 2020)
@@
SAR (Bui et al., 2019)
HaPiM (Zhou et al., 2023b)
[Cloze Test
@@
JavaBERT (de Sousa & Hasselbring, 2021)
@@
Repair
@@
Refactory (Hu et al., 2019), PyTER (Oh & Oh, 2022)
@@
DEAR (Li et al., 2022f)
@@
Conversational APR (Xia & Zhang, 2023), Jiang et al. (2023a),
@@
TypeFix (Peng et al., 2023b)
@@
Infilling
@@
StarCoder (Li et al., 2023g), Code LLaMA (Rozière et al., 2023)
[Code-to-Text]
[Code-to-Pattern]
[Text-to-Text]
@@
where level=2text width=3.5em,font=,
@@
[Code Evaluation
[Text-to-Code
[Code-to-Code
[Deobfuscation
@@
Debin (He et al., 2018), JSNeat (Tran et al., 2019)
DIRE (Lacomis et al., 2019), Artuso et al. (2021)
@@
SLaDe (Armengol-Estapé et al., 2023), LmPa (Xu et al., 2023b)
[Unit Test Generation
EvoSuite (Fraser & Arcuri, 2011), EvoSuiteR (Shamshiri, 2015),
TSE (Shimmi & Rahimi, 2022), Nessie (Arteca et al., 2022)
AthenaTest (Tufano et al., 2021a), FSLM (Bareiß et al., 2022),
TestPilot (Schäfer et al., 2023), A3Test (Alagarsamy et al., 2023),
TeCo (Nie et al., 2023), CodaMosa (Lemieux et al., 2023),
ChatTester (Yuan et al., 2023b), ChatUniTest (Xie et al., 2023b)
[Assertion Generation
MeMo (Blasi et al., 2021)
Atlas (Watson et al., 2020)
@@
FSLM (Bareiß et al., 2022), Yuan et al. (2023a)
@@
Generation
@@
IBIR (Khanfir et al., 2023b), PBMT (Bartocci et al., 2023)
@@
SemSeed (Patra & Pradel, 2021)
@@
FSLM (Bareiß et al., 2022)
[Fuzzing
@@
Fuzz (Yang et al., 2023c)
@@
PreFuzz (Wu et al., 2022)
@@
WhiteFox (Yang et al., 2023b)
[Code-to-Text]
[Code-to-Pattern]
[Text-to-Text]
@@
- Code repair, also known as bug fix, aims to fix a piece of buggy code. Like code translation, it is a traditional sequence-to-sequence generation task, and surveys are abundant on this topic (Gazzola et al., 2018; Monperrus, 2018; Zhong et al., 2022; Zhang et al., 2023c; Huang et al., 2023a). A generalization of code repair is code edit or code change, where the task is to model and predict code changes (such as Git commits), which are not necessarily about bug fixing.
- Code infilling is another recently proposed task, after fill-in-the-middle pretraining (Bavarian et al., 2022) became popular. It is a generalization of code completion, where not only the left context, but also the right context is given. However, it differs from cloze test in that the target of cloze test is only one token, while the target of code infilling can be an entire line or even multiple lines, which requires a decoder to generate autoregressively.
@@
- Unit test generation aims to generate unit tests for a given program. Prior to the rise of Codex and other code LLMs, almost all works in this area employed non-neural methods (see Figure 6). In the age of LLMs, however, this task is ever more important, as researches have shown that the current unit tests for evaluating LLMs’ program synthesis capability may be insufficient (Liu et al., 2023f).
@@
[Code Evaluation
[Text-to-Code]
[Code-to-Code]
[Code-to-Text
[Code Summarization
CloCom (Wong et al., 2015)
@@
MLCS (Zhou et al., 2023c)
TranS (Wang et al., 2020e), Ahmad et al. (2020),
@@
Su & McMillan (2023b)
@@
Review
@@
ClusterChanges (Barnett et al., 2015), CommentFinder (Hong et al., 2022)
@@
DACE (Shi et al., 2019), CORE (Siow et al., 2020)
@@
LLaMA-Reviewer (Lu et al., 2023)
[Identifier Prediction
@@
HeMa (Jiang et al., 2019)
@@
DMACOS (Xie et al., 2021), NamPat (Yang et al., 2022)
@@
Mario (Wang et al., 2023b)
[Commit Message Generation
@@
NNGen (Liu et al., 2018), ChangeDoc (Huang et al., 2020)
@@
CC2Vec (Hoang et al., 2020), CoRec (Wang et al., 2021b)
@@
COME (He et al., 2023a), ExGroFi (Wang et al., 2023a)
[Code-to-Pattern
[Text-to-Text
@@
where level=2text width=3.5em,font=,
@@
[Code Evaluation
[Text-to-Code]
[Code-to-Code]
[Code-to-Text]
[Code-to-Pattern
@@
Prediction
@@
Typpete (Hassan et al., 2018)
@@
PYInfer (Cui et al., 2021)
@@
TypeT5 (Wei et al., 2023), OpenTau (Cassano et al., 2023b),
TypeGen (Peng et al., 2023c)
@@
Detection
@@
NAR-Miner (Bian et al., 2018), (Pearce et al., 2022)
Wang et al. (2016b), VulDeePecker (Li et al., 2018c),
@@
ReGVD (Nguyen et al., 2022)
@@
Gao et al. (2023d)
@@
Detection
@@
NIL (Nakagawa et al., 2021)
@@
Fang et al. (2020), Yahya & Kim (2022)
@@
Reasoning
@@
CMMLU (Li et al., 2023a),
CodeApex (Fu et al., 2023)
[Code/Author Classification
SCAP (Frantzeskou et al., 2011), Mahbub et al. (2022)
@@
ASTNN (Zhang et al., 2019a), InferCode (Bui et al., 2021a)
Puri et al. (2021), Wang et al. (2022a)
@@
Detection
DetectGPT4Code (Yang et al., 2023e)
[Text-to-Text
[Document Translation
[Lu et al. (2021)]
@@
Analysis
@@
Logram (Dai et al., 2022), SPINE (Wang et al., 2022e)
@@
UniParser (Liu et al., 2022e)
@@
LogPrompt (Liu et al., 2023k), LogGPT (Qi et al., 2023),
Mudgal & Wouhaybi (2023), LLMParser (Jiang et al., 2023b)

### v5

@@ 2.1.1 Text-to-Code
Text-to-code tasks take text as input, and output code. Related works are listed in Figure 3
@@ 2.1.2 Code-to-Code
[Code Evaluation, fill=black, fill opacity=0.7, text=white
[Text-to-Code, fill=yellow, fill opacity=0.5
[Code-to-Code, fill=cyan, fill opacity=0.2
[Code Search, fill=cyan, fill opacity=0.2
@@
COSAL (Mathew & Stolee, 2021), fill=cyan, fill opacity=0.2
[, fill=cyan, fill opacity=0.2
Corder (Bui et al., 2021b), Cosco (Saieva et al., 2023), fill=cyan, fill opacity=0.2
@@
Completion, fill=cyan, fill opacity=0.2
@@
Deep3 (Raychev et al., 2016a), Hellendoorn & Devanbu (2017), fill=cyan, fill opacity=0.2
@@
SLM (Alon et al., 2020), Pythia (Svyatkovskiy et al., 2019), NLM (Karampatsis et al., 2020), fill=cyan, fill opacity=0.2
@@
LongCoder (Guo et al., 2023b), fill=cyan, fill opacity=0.2
@@
Translation, fill=cyan, fill opacity=0.2
@@
mppSMT (Nguyen et al., 2015), fill=cyan, fill opacity=0.2
Tree2Tree (Chen et al., 2018), Grammar Tree2Tree (Drissi et al., 2018), fill=cyan, fill opacity=0.2
@@
Self-Debugging (Chen et al., 2023d),
@@
Pan et al. (2023),
Explain-then-Translate (Tang et al., 2023b),
Macedo et al. (2024), fill=cyan, fill opacity=0.2
@@
Mining, fill=cyan, fill opacity=0.2
RASH (Zhang et al., 2018), RAPIM (Alrubaye et al., 2020), fill=cyan, fill opacity=0.2
@@
SAR (Bui et al., 2019), fill=cyan, fill opacity=0.2
HaPiM (Zhou et al., 2023b), fill=cyan, fill opacity=0.2
[Cloze Test, fill=cyan, fill opacity=0.2
[, fill=cyan, fill opacity=0.2
[, fill=cyan, fill opacity=0.2
@@
JavaBERT (de Sousa & Hasselbring, 2021), fill=cyan, fill opacity=0.2
@@
Repair, fill=cyan, fill opacity=0.2
@@
Refactory (Hu et al., 2019), PyTER (Oh & Oh, 2022), fill=cyan, fill opacity=0.2
@@
DEAR (Li et al., 2022f), fill=cyan, fill opacity=0.2
@@
Conversational APR (Xia & Zhang, 2023), Jiang et al. (2023b),
@@
TypeFix (Peng et al., 2023b),
Maniple (Parasaram et al., 2024), fill=cyan, fill opacity=0.2
@@
Infilling, fill=cyan, fill opacity=0.2
[, fill=cyan, fill opacity=0.2
[, fill=cyan, fill opacity=0.2
@@
StarCoder (Li et al., 2023h), Code LLaMA (Rozière et al., 2023), fill=cyan, fill opacity=0.2
[Code-to-Text, fill=Orange, fill opacity=0.2]
[Code-to-Pattern, fill=VioletRed, fill opacity=0.2]
[Text-to-Text, fill=Dandelion, fill opacity=0.4]
@@
where level=2text width=4.1em,font=,
@@
[Code Evaluation, fill=black, fill opacity=0.7, text=white
[Text-to-Code, fill=yellow, fill opacity=0.5
[Code-to-Code, fill=cyan, fill opacity=0.2
[Decompilation, fill=cyan, fill opacity=0.2
, fill=cyan, fill opacity=0.2
Neutron (Liang et al., 2021),
NeurDP (Cao et al., 2022)
, fill=cyan, fill opacity=0.2
BTC (Hosseini & Dolan-Gavitt, 2022),
SLaDe (Armengol-Estapé et al., 2023),
Nova+ (Jiang et al., 2023c),
LLM4Decompile (Tan et al., 2024)
, fill=cyan, fill opacity=0.2
[Deobfuscation, fill=cyan, fill opacity=0.2
@@
Debin (He et al., 2018), JSNeat (Tran et al., 2019), fill=cyan, fill opacity=0.2
DIRE (Lacomis et al., 2019), Artuso et al. (2021), fill=cyan, fill opacity=0.2
@@
LmPa (Xu et al., 2023), fill=cyan, fill opacity=0.2
[Assertion Generation, fill=cyan, fill opacity=0.2
MeMo (Blasi et al., 2021), fill=cyan, fill opacity=0.2
Atlas (Watson et al., 2020), fill=cyan, fill opacity=0.2
@@
FSLM (Bareiß et al., 2022), Yuan et al. (2023a), fill=cyan, fill opacity=0.2
@@
Generation, fill=cyan, fill opacity=0.2
@@
IBIR (Khanfir et al., 2023b), PBMT (Bartocci et al., 2023), fill=cyan, fill opacity=0.2
@@
SemSeed (Patra & Pradel, 2021), fill=cyan, fill opacity=0.2
@@
FSLM (Bareiß et al., 2022), fill=cyan, fill opacity=0.2
[Unit Test Generation, fill=cyan, fill opacity=0.2
EvoSuite (Fraser & Arcuri, 2011), EvoSuiteR (Shamshiri, 2015),
TSE (Shimmi & Rahimi, 2022), Nessie (Arteca et al., 2022), fill=cyan, fill opacity=0.2
[, fill=cyan, fill opacity=0.2
AthenaTest (Tufano et al., 2021a), FSLM (Bareiß et al., 2022),
TestPilot (Schäfer et al., 2023), A3Test (Alagarsamy et al., 2023),
TeCo (Nie et al., 2023), CodaMosa (Lemieux et al., 2023),
ChatTester (Yuan et al., 2023b), ChatUniTest (Xie et al., 2023b),
Tang et al. (2023a),
PBT-GPT (Vikram et al., 2023),
Shin et al. (2023),
MuTAP (Dakhel et al., 2023),
Zhang et al. (2023f),
RLSQM (Steenhoek et al., 2023b),
Guilherme & Vincenzi (2023),
CoverUp (Pizzorno & Berger, 2024),
TELPA (Yang et al., 2024)
, fill=cyan, fill opacity=0.2
[Fuzzing, fill=cyan, fill opacity=0.2
@@
Fuzz (Yang et al., 2023c), fill=cyan, fill opacity=0.2
@@
PreFuzz (Wu et al., 2022), fill=cyan, fill opacity=0.2
@@
Fuzz4All (Xia et al., 2023a),
WhiteFox (Yang et al., 2023b), fill=cyan, fill opacity=0.2
[Code-to-Text, fill=Orange, fill opacity=0.2]
[Code-to-Pattern, fill=VioletRed, fill opacity=0.2]
[Text-to-Text, fill=Dandelion, fill opacity=0.4]
@@
- Code repair, also known as bug fix, aims to fix a piece of buggy code. Like code translation, it is a traditional sequence-to-sequence generation task, and surveys are abundant on this topic (Gazzola et al., 2018; Monperrus, 2018; Zhong et al., 2022; Zhang et al., 2023d; Huang et al., 2023a). A generalization of code repair is code edit or code change, where the task is to model and predict code changes (such as Git commits), which are not necessarily about bug fixing.
- Code infilling is another recently proposed task, after fill-in-the-middle pretraining (Bavarian et al., 2022) became popular. It is a generalization of code completion, where not only the left context, but also the right context is given. However, it differs from cloze test in that the target of cloze test is only one token, while the target of code infilling can be an entire line or even multiple lines, which requires a decoder to generate autoregressive.
- Decompilation is the reverse process of compilation. In this task a model takes a compiled program - represented in assembly code or binary machine code - and aims to output the original source program in high-level languages such as C.
@@
- Unit test generation aims to generate unit tests for a given program. Prior to the rise of Codex and other code LLMs, almost all works in this area employed non-neural methods (see Figure 5). In the age of LLMs, however, this task is ever more important, as research has shown that the current unit tests for evaluating LLMs’ program synthesis capability may be insufficient (Liu et al., 2023f). Wang et al. (2023a) provide a comprehensive survey on software testing with LLMs.
@@
[Code Evaluation, fill=black, fill opacity=0.7, text=white
[Text-to-Code, fill=yellow, fill opacity=0.5]
[Code-to-Code, fill=cyan, fill opacity=0.4]
[Code-to-Text, fill=Orange, fill opacity=0.2
[Code Summarization, fill=Orange, fill opacity=0.2
CloCom (Wong et al., 2015), fill=Orange, fill opacity=0.2
@@
MLCS (Zhou et al., 2023c), fill=Orange, fill opacity=0.2
TranS3 (Wang et al., 2020e), Ahmad et al. (2020),
@@
Su & McMillan (2023b),
CSA-Trans (Oh & Yoo, 2024),
Haldar & Hockenmaier (2024)
, fill=Orange, fill opacity=0.2
@@
Review, fill=Orange, fill opacity=0.2
@@
ClusterChanges (Barnett et al., 2015), CommentFinder (Hong et al., 2022), fill=Orange, fill opacity=0.2
@@
DACE (Shi et al., 2019), CORE (Siow et al., 2020), fill=Orange, fill opacity=0.2
@@
LLaMA-Reviewer (Lu et al., 2023a), fill=Orange, fill opacity=0.2
[Identifier Prediction, fill=Orange, fill opacity=0.2
@@
HeMa (Jiang et al., 2019), fill=Orange, fill opacity=0.2
@@
DMACOS (Xie et al., 2021), NamPat (Yang et al., 2022), fill=Orange, fill opacity=0.2
@@
Mario (Wang et al., 2023c), fill=Orange, fill opacity=0.2
[Commit Message Generation, fill=Orange, fill opacity=0.2
@@
NNGen (Liu et al., 2018), ChangeDoc (Huang et al., 2020), fill=Orange, fill opacity=0.2
@@
CC2Vec (Hoang et al., 2020), CoRec (Wang et al., 2021b), fill=Orange, fill opacity=0.2
@@
COME (He et al., 2023a), ExGroFi (Wang et al., 2023b), fill=Orange, fill opacity=0.2
[Code-to-Pattern, fill=VioletRed, fill opacity=0.2
[Text-to-Text, fill=Dandelion, fill opacity=0.4
@@
where level=2text width=4.1em,font=,
@@
[Code Evaluation, fill=black, fill opacity=0.7, text=white
[Text-to-Code, fill=yellow, fill opacity=0.5]
[Code-to-Code, fill=cyan, fill opacity=0.4]
[Code-to-Text, fill=Orange, fill opacity=0.2]
[Code-to-Pattern, fill=VioletRed, fill opacity=0.2
@@
Prediction, fill=VioletRed, fill opacity=0.2
@@
Typpete (Hassan et al., 2018), fill=VioletRed, fill opacity=0.2
@@
PYInfer (Cui et al., 2021), fill=VioletRed, fill opacity=0.2
@@
TypeT5 (Wei et al., 2023), OpenTau (Cassano et al., 2023c),
TypeGen (Peng et al., 2023c),
Lucchetti & Guha (2024)
, fill=VioletRed, fill opacity=0.2
@@
Detection, fill=VioletRed, fill opacity=0.2
@@
NAR-Miner (Bian et al., 2018), (Pearce et al., 2022), fill=VioletRed, fill opacity=0.2
Wang et al. (2016b), VulDeePecker (Li et al., 2018b),
@@
ReGVD (Nguyen et al., 2022), fill=VioletRed, fill opacity=0.2
@@
Gao et al. (2023d),
Ullah et al. (2023),
LLM4Vuln (Sun et al., 2024),
MuCoLD (Mao et al., 2024),
Steenhoek et al. (2024)
, fill=VioletRed, fill opacity=0.2
[Malware
Detection, fill=VioletRed, fill opacity=0.2
, fill=VioletRed, fill opacity=0.2
, fill=VioletRed, fill opacity=0.2
Malbert (Xu et al., 2021b),
SHERLOCK (Seneviratne et al., 2022),
MSDT (Tsfaty & Fire, 2022),
MalBERTv2 (Rahali & Akhloufi, 2023),
SocketAI Scanner (Zahan et al., 2024)
, fill=VioletRed, fill opacity=0.2
@@
Detection, fill=VioletRed, fill opacity=0.2
@@
NIL (Nakagawa et al., 2021), fill=VioletRed, fill opacity=0.2
@@
Fang et al. (2020), Yahya & Kim (2022), fill=VioletRed, fill opacity=0.2
@@
Reasoning, fill=VioletRed, fill opacity=0.2
[, fill=VioletRed, fill opacity=0.2
[, fill=VioletRed, fill opacity=0.2
@@
CMMLU (Li et al., 2023b),
CodeApex (Fu et al., 2023), fill=VioletRed, fill opacity=0.2
[Code/Author Classification, fill=VioletRed, fill opacity=0.2
SCAP (Frantzeskou et al., 2011), Mahbub et al. (2022), fill=VioletRed, fill opacity=0.2
@@
ASTNN (Zhang et al., 2019a), InferCode (Bui et al., 2021a), fill=VioletRed, fill opacity=0.2
Puri et al. (2021), Wang et al. (2022a), fill=VioletRed, fill opacity=0.2
@@
Detection, fill=VioletRed, fill opacity=0.2
[, fill=VioletRed, fill opacity=0.2
[, fill=VioletRed, fill opacity=0.2
DetectGPT4Code (Yang et al., 2023e), fill=VioletRed, fill opacity=0.2
[Text-to-Text, fill=Dandelion, fill opacity=0.4
[Document Translation, fill=Dandelion, fill opacity=0.4
[, fill=Dandelion, fill opacity=0.4
[, fill=Dandelion, fill opacity=0.4
[Lu et al. (2021), fill=Dandelion, fill opacity=0.4]
@@
Analysis, fill=Dandelion, fill opacity=0.4
@@
Logram (Dai et al., 2022), SPINE (Wang et al., 2022e), fill=Dandelion, fill opacity=0.4
@@
UniParser (Liu et al., 2022e), fill=Dandelion, fill opacity=0.4
@@
LogPrompt (Liu et al., 2023l), LogGPT (Qi et al., 2023),
Mudgal & Wouhaybi (2023), LLMParser (Jiang et al., 2023d), fill=Dandelion, fill opacity=0.4